OpenStudy (amberofficial1):

Help Immediately!! if a 5000N force is applied to a 10,000kg object traveling at 10 m/s for 15 seconds, does the object come to a complete stop? have number to prove it please

OpenStudy (ytrewqmiswi):

force=mass x acceleration 5000=10000 x acceleration 5000/10000=acceleration 1/2=acceleration v=u+at where v=final velocity u=initial velocity a=acceleration t=time lets find the time 't' when the body stops so because the body stops v=0 and u is given as 10m/s and a is 1/2 we will take it as -1/2 cause the force is opposing the motion so we have- v=u+at 0=10-1/2t -10=-1/2t t=20secs so the body doesn't stop in 15secs rather it stops after 20secs

OpenStudy (amberofficial1):

thanks!!

OpenStudy (ytrewqmiswi):

np